OpenStudy (mysticfury):

Variable y varies directly with variable x, and y = 6 when x = 4. Enter the constant of variation in lowest terms.

OpenStudy (mysticfury):

 you have to solve for the question marks

OpenStudy (mysticfury):

@Directrix

OpenStudy (mysticfury):

@pooja195 @Daniee_Bruhh

OpenStudy (mysticfury):

@retirEEd

OpenStudy (daniee_bruhh):

@Dani_Rose @Directrix

OpenStudy (universetoast):

@MysticFury is it asking you to substitute?

OpenStudy (mysticfury):

idk @UniverseToast

OpenStudy (mysticfury):

thats the whole question

OpenStudy (universetoast):

Can you send me a picture of the whole question?

Directrix (directrix):

Direct variation means that the equation will be of the form: y = k*x where k is the constant of variation Replace y be 6 and x by 4 and solve the equation for k.

OpenStudy (universetoast):

OH, yeah @Directrix is right. Oops

OpenStudy (mysticfury):

so what 1.5? it needs to be a fraction

Directrix (directrix):

y = k*x 6 = k* 4 Solve for k @MysticFury

Directrix (directrix):

@MysticFury What did you get for the answer?

OpenStudy (mysticfury):

i got 1.5

Directrix (directrix):

You are to enter it as a fraction according to the posted question.

Directrix (directrix):

Look at the screen shot.

OpenStudy (mysticfury):

lots of fractions equal 1.5

Directrix (directrix):

1.5 = ? as a fraction.

OpenStudy (mysticfury):

3/2?

Directrix (directrix):

True. But, the question asks for simplest terms. That narrows the fraction to exactly one.

Directrix (directrix):

>> 3/2? Correct.

OpenStudy (mysticfury):

ok thank you

Directrix (directrix):

You are welcome.

OpenStudy (needhelpwithk12):

what grade are you in first of all?